Becky’s Bikes Inc. has two divisions: Retail and Service. The following information is for each division at Becky’s Bikes for the most recent fiscal year. Retail Division Service Division Income Statement Sales $1,200,000 $1,000,000 Cost of sales 550,000 400,000 Gross margin 650,000 600,000 Allocated overhead (from corporate) 210,000 180,000 Marketing expense 130,000 120,000 Administrative expense 60,000 55,000 Operating income 250,000 245,000 Income tax expense (40% rate) 100,000 98,000 Net income 150,000 147,000 Balance Sheet Information Average operating assets $600,000 $375,000 Non-interest bearing current liabilities 100,000 75,000 Percent cost of capital 12% 12% To calculate EVA, management requires adjustments for marketing and non-interest bearing current liabilities as outlined below. Marketing will be capitalized and amortized over several years resulting in an increase to average operating assets of $45,000 for the Retail division and $37,500 for the Services division. On the income statement, marketing expense for the year will be added back to operating income, then marketing amortization expense for one year will be deducted. The current year amortization expense will total $30,000 for the Retail division and $20,000 for the Services division. Non-interest bearing liabilities will be deducted from average operating assets. Required Calculate economic value added (EVA) for each division and comment on your results.
Answer #1

Answers Retail Division net operating profit afater taxes = (1,50,000 + (1,30,000x0.60) -(30,000 X 0.6) = 150,000+ (72000)– (:. Econoncie volue added = 464,500 005991 - - 207,000 -- 40500 Economic value added = 207,000-337,500 x 124, Qoste 000St-Oe